OpenStudy (anonymous):

I really suck at dividing equations. Multiplication Equations: 10x=50

OpenStudy (anonymous):

10/50=5, so x =5

OpenStudy (anonymous):

10x=50 x=50/10 x=5

OpenStudy (anonymous):

moon can u help me on my other 1

jhonyy9 (jhonyy9):

comen where is ?

OpenStudy (anonymous):

srry rowerspike wasnt there......................